# Blind Inverse Game Theory: Jointly Decoding Rewards and Rationality in Entropy-Regularized Competitive Games

### v1

@@ Appendix B Background and Derivations
B.1 Derivation of the Linearized QRE Constraints
We detail the derivation of the linearized QRE constraints (Eq. 4) from the QRE fixed-point equations (Eq. 2 and 3).
Recall the QRE definition for Player 1 (Eq. 2):
$\mu^{*}(a)=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(a^{\prime},\cdot)\nu^{*}/\tau)}.$ (17)
@@ B.1 Derivation of the Linearized QRE Constraints
$\frac{\mu^{*}(a)}{\mu^{*}(1)}=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)/Z}{\exp(Q(1,\cdot)\nu^{*}/\tau)/Z}=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\exp(Q(1,\cdot)\nu^{*}/\tau)},$ (18)
@@
$\log\left(\frac{\mu^{*}(a)}{\mu^{*}(1)}\right)$ $=\log(\exp(Q(a,\cdot)\nu^{*}/\tau))-\log(\exp(Q(1,\cdot)\nu^{*}/\tau))$ (19)
$=\frac{1}{\tau}(Q(a,\cdot)\nu^{*}-Q(1,\cdot)\nu^{*})$ (20)
$=\frac{1}{\tau}(Q(a,\cdot)-Q(1,\cdot))\nu^{*}.$ (21)
Rearranging the terms yields the first set of linear constraints in Eq. 4:
$(Q(a,\cdot)-Q(1,\cdot))\nu^{*}=\tau\cdot\log\left(\frac{\mu^{*}(a)}{\mu^{*}(1)}\right),\quad\forall a\in\mathcal{A}\setminus\{1\}.$ (22)
Similarly, for Player 2 (Eq. 3), we analyze the ratio of $\nu^{*}(b)$ to $\nu^{*}(1)$:
$\frac{\nu^{*}(b)}{\nu^{*}(1)}=\frac{\exp(-Q(\cdot,b)^{\top}\mu^{*}/\tau)}{\exp(-Q(\cdot,1)^{\top}\mu^{*}/\tau)}.$ (23)
@@
$\log\left(\frac{\nu^{*}(b)}{\nu^{*}(1)}\right)$ $=\frac{1}{\tau}(-Q(\cdot,b)^{\top}\mu^{*}-(-Q(\cdot,1)^{\top}\mu^{*}))$ (24)
$=\frac{1}{\tau}(Q(\cdot,1)^{\top}-Q(\cdot,b)^{\top})\mu^{*}$ (25)
$=\frac{1}{\tau}(Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}.$ (26)
Rearranging the terms yields the second set of linear constraints in Eq. 4:
$(Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}=\tau\cdot\log\left(\frac{\nu^{*}(b)}{\nu^{*}(1)}\right),\quad\forall b\in\mathcal{B}\setminus\{1\}.$ (27)
@@
B.2 Key Definitions and Theorems from Matrix Analysis
@@
Definition 1 (Matrix Norms).
@@
Definition 2 (Moore-Penrose Pseudoinverse).
@@
Theorem 1 (Weyl’s Inequality for Singular Values).
@@
This theorem is crucial for establishing the stability of the matrix $X$ under small perturbations (Step 1 in the proof of Theorem 2).
B.3 Key Theorems from Concentration Inequalities
@@
Theorem 2 (Bretagnolle-Huber-Carol Inequality).
@@
$P(\|\hat{\mu}-\mu^{*}\|_{1}>\epsilon)\leq 2^{m}e^{-N\epsilon^{2}/2}.$ (29)
This theorem provides the finite-sample bound on the L1 distance between the estimated and true policies (Lemma 1).
Appendix C Proofs for Matrix Games
@@
C.1 Proof of Theorem 1 (Identifiability of Blind-IGT)
@@
Proof of Sufficiency (If): Assume the Rank Condition ($\text{rank}(X^{*})=d$) and the Non-Uniformity Condition ($y^{*}\neq 0$) hold.
@@
$\theta^{*}=(X^{*})^{\dagger}(\tau^{*}y^{*})=\tau^{*}(X^{*})^{\dagger}y^{*}.$ (30)
@@
$\|\theta^{*}\|_{2}=\|\tau^{*}\theta_{LS}^{*}\|_{2}=|\tau^{*}|\|\theta_{LS}^{*}\|_{2}=1.$ (31)
@@
$\theta^{*}=\tau^{*}\theta_{LS}^{*}=\frac{\theta_{LS}^{*}}{\|\theta_{LS}^{*}\|_{2}}.$ (33)
Proof of Necessity (Only If): We prove the contrapositive by considering the failure of each condition.
@@
C.2 Supporting Lemmas for Finite Sample Analysis
We present and prove the supporting lemmas required for the proof of Theorem 2. These lemmas characterize the concentration of the policy estimates and the resulting perturbations in the system matrices.
Lemma 1 (Policy Estimation Error).
@@ C.2 Supporting Lemmas for Finite Sample Analysis
By the Bretagnolle-Huber-Carol inequality (Theorem 2 in Appendix B), for any $\epsilon>0$:
$P(\|\hat{\mu}-\mu^{*}\|_{1}>\epsilon)\leq 2^{m}e^{-N\epsilon^{2}/2}.$ (34)
@@
$\frac{\delta}{2}=2^{m}e^{-N\epsilon_{\mu}^{2}/2}\implies\log(2\cdot 2^{m}/\delta)=N\epsilon_{\mu}^{2}/2\implies\epsilon_{\mu}=\sqrt{\frac{2\log(2\cdot 2^{m}/\delta)}{N}}.$ (35)
@@
Lemma 2 (Perturbation of X and y).
Let $L$ be the bound on $\|\phi\|_{2}$ (Assumption 1). Under Assumption 3, let $\xi>0$ be the minimum probability. If $N$ is large enough such that the policy estimation errors satisfy $\epsilon_{\mu}<\xi/2$ and $\epsilon_{\nu}<\xi/2$, then with high probability (conditioned on the event of Lemma 1):
@@
$\|(A(\hat{\nu})-A(\nu^{*}))_{a-1}\|_{2}$ $=\left\|\sum_{b^{\prime}\in\mathcal{B}}(\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime}))(\phi(a,b^{\prime})-\phi(1,b^{\prime}))^{\top}\right\|_{2}$ (38)
$\leq\sum_{b^{\prime}\in\mathcal{B}}|\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime})|\cdot\|\phi(a,b^{\prime})-\phi(1,b^{\prime})\|_{2}.$ (39)
@@
$\|(A(\hat{\nu})-A(\nu^{*}))_{a-1}\|_{2}$ $\leq 2L\sum_{b^{\prime}\in\mathcal{B}}|\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime})|=2L\|\hat{\nu}-\nu^{*}\|_{1}\leq 2L\epsilon_{\nu}.$ (40)
@@
$\|A(\hat{\nu})-A(\nu^{*})\|_{F}\leq\sqrt{(m-1)(2L\epsilon_{\nu})^{2}}\leq 2L\sqrt{m}\epsilon_{\nu}.$ (41)
@@
$\|\hat{X}-X^{*}\|_{op}$ $\leq\|\hat{X}-X^{*}\|_{F}=\sqrt{\|A(\hat{\nu})-A(\nu^{*})\|_{F}^{2}+\|B(\hat{\mu})-B(\mu^{*})\|_{F}^{2}}$ (42)
$\leq\sqrt{4L^{2}m\epsilon_{\nu}^{2}+4L^{2}n\epsilon_{\mu}^{2}}=2L\sqrt{m\epsilon_{\nu}^{2}+n\epsilon_{\mu}^{2}}.$ (43)
@@
$\|c(\hat{\mu})-c(\mu^{*})\|_{2}^{2}$ $=\sum_{a=2}^{m}|c(\hat{\mu})_{a-1}-c(\mu^{*})_{a-1}|^{2}$ (49)
$\leq\sum_{a=2}^{m}\left(\frac{2}{\xi}\right)^{2}(|\hat{\mu}(a)-\mu^{*}(a)|+|\hat{\mu}(1)-\mu^{*}(1)|)^{2}$ (50)
$\leq\frac{8}{\xi^{2}}\sum_{a=2}^{m}(|\hat{\mu}(a)-\mu^{*}(a)|^{2}+|\hat{\mu}(1)-\mu^{*}(1)|^{2})\quad(\text{Since }(x+y)^{2}\leq 2(x^{2}+y^{2}))$ (51)
$=\frac{8}{\xi^{2}}\left(\sum_{a=2}^{m}|\hat{\mu}(a)-\mu^{*}(a)|^{2}+(m-1)|\hat{\mu}(1)-\mu^{*}(1)|^{2}\right)$ (52)
$\leq\frac{8m}{\xi^{2}}\|\hat{\mu}-\mu^{*}\|_{2}^{2}.$ (53)
@@
$\|c(\hat{\mu})-c(\mu^{*})\|_{2}\leq\frac{\sqrt{8m}}{\xi}\epsilon_{\mu}.$ (54)
@@
$\|\hat{y}-y^{*}\|_{2}$ $=\sqrt{\|c(\hat{\mu})-c(\mu^{*})\|_{2}^{2}+\|d(\hat{\nu})-d(\nu^{*})\|_{2}^{2}}$ (55)
$\leq\sqrt{\frac{8m}{\xi^{2}}\epsilon_{\mu}^{2}+\frac{8n}{\xi^{2}}\epsilon_{\nu}^{2}}=\frac{\sqrt{8}}{\xi}\sqrt{m\epsilon_{\mu}^{2}+n\epsilon_{\nu}^{2}}.$ (56)
@@
Lemma 3 (Stability of Normalization).
@@
$\left\|\frac{v_{1}}{\|v_{1}\|_{2}}-\frac{v_{0}}{\|v_{0}\|_{2}}\right\|_{2}=\left\|\frac{v_{1}\|v_{0}\|_{2}-v_{0}\|v_{1}\|_{2}}{\|v_{1}\|_{2}\|v_{0}\|_{2}}\right\|_{2}.$ (57)
@@
$\|v_{1}\|v_{0}\|_{2}-v_{0}\|v_{1}\|_{2}\|$ $=\|(v_{0}+\Delta v)\|v_{0}\|_{2}-v_{0}(\|v_{0}\|_{2}+\Delta_{\|v\|})\|$ (58)
$=\|v_{0}\|v_{0}\|_{2}+\Delta v\|v_{0}\|_{2}-v_{0}\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|$ (59)
$=\|\Delta v\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|.$ (60)
@@
$\|\Delta v\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|$ $\leq\|\Delta v\|v_{0}\|_{2}\|+\|v_{0}\Delta_{\|v\|}\|$ (61)
$=\|\Delta v\|_{2}\|v_{0}\|_{2}+|\Delta_{\|v\|}|\|v_{0}\|_{2}$ (62)
$\leq\epsilon\|v_{0}\|_{2}+\epsilon\|v_{0}\|_{2}=2\epsilon\|v_{0}\|_{2}.$ (63)
@@
$\left\|\frac{v_{1}}{\|v_{1}\|_{2}}-\frac{v_{0}}{\|v_{0}\|_{2}}\right\|_{2}\leq\frac{2\epsilon\|v_{0}\|_{2}}{(\delta/2)\|v_{0}\|_{2}}=\frac{4\epsilon}{\delta}.$ (65)
C.3 Proof of Theorem 2 (Finite Sample Bounds)
@@
Let $\sigma^{*}=\sigma_{\min}(X^{*})$. Since the Rank Condition holds (Theorem 1), $\sigma^{*}>0$. Let $\epsilon_{N}$ be the policy estimation error rate from Lemma 1. Let $C_{X},C_{Y}$ be the perturbation constants from Lemma 2.
@@
Stability of the Estimated System. We need to ensure that the estimated matrix $\hat{X}$ remains full rank and well-conditioned, provided $N$ is sufficiently large. We use Weyl’s inequality (Theorem 1 in Appendix B) for the perturbation of singular values:
$\sigma_{\min}(\hat{X})\geq\sigma_{\min}(X^{*})-\|\hat{X}-X^{*}\|_{op}=\sigma^{*}-\|\hat{X}-X^{*}\|_{op}.$ (66)
We require $\|\hat{X}-X^{*}\|_{op}<\sigma^{*}/2$ to guarantee $\sigma_{\min}(\hat{X})>\sigma^{*}/2$. By Lemma 2, this holds if $C_{X}\epsilon_{N}<\sigma^{*}/2$. This defines the primary condition for ”sufficiently large NN”. We also require $N$ large enough such that the conditions of Lemma 2 (i.e., $\epsilon_{\mu},\epsilon_{\nu}<\xi/2$) hold.
@@
We utilize standard perturbation bounds for the least squares solution when the true system is consistent (zero residual). Since $\hat{X}$ is full rank (Step 1), $\hat{\theta}_{LS}=\hat{X}^{\dagger}\hat{y}$. The error bound is given by (e.g., Theorem 20.1 in Higham2002, specialized for zero residual):
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}\leq\|\hat{X}^{\dagger}\|_{2}(\|\hat{X}-X^{*}\|_{op}\|\theta^{\prime*}\|_{2}+\|\hat{y}-y^{*}\|_{2})+\mathcal{O}(\epsilon_{N}^{2}).$ (67)
Since $\sigma_{\min}(\hat{X})\geq\sigma^{*}/2$, we have $\|\hat{X}^{\dagger}\|_{2}=1/\sigma_{\min}(\hat{X})\leq 2/\sigma^{*}$. We substitute the bounds from Lemma 2 and note that $\|\theta^{\prime*}\|_{2}=\|\theta^{*}\|_{2}/\tau^{*}=1/\tau^{*}$ (since $C=1$).
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}$ $\leq\frac{2}{\sigma^{*}}\left((C_{X}\epsilon_{N})\frac{1}{\tau^{*}}+C_{Y}\epsilon_{N}\right)$ (68)
$=\left(\frac{2C_{X}}{\sigma^{*}\tau^{*}}+\frac{2C_{Y}}{\sigma^{*}}\right)\epsilon_{N}=:C_{LS}\epsilon_{N}.$ (69)
@@
Error in $\hat{\theta}$. We apply Lemma 3 (Stability of Normalization) with $v_{1}=\hat{\theta}_{LS}$ and $v_{0}=\theta^{\prime*}$. The error is $\epsilon=\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}\leq C_{LS}\epsilon_{N}$. The norm is $\delta=\|\theta^{\prime*}\|_{2}=1/\tau^{*}$. We require $\epsilon<\delta/2$, i.e., $C_{LS}\epsilon_{N}<1/(2\tau^{*})$. This holds for sufficiently large $N$ since $\epsilon_{N}\rightarrow 0$.

### v2

@@ Appendix C Background and Derivations
C.1 Derivation of the Linearized QRE Constraints
We detail the derivation of the linearized QRE constraints (Eq. 3.4) from the QRE fixed-point equations (Eq. 3.2 and 3.3).
Recall the QRE definition for Player 1 (Eq. 3.2):
$\mu^{*}(a)=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(a^{\prime},\cdot)\nu^{*}/\tau)}.$ (C.1)
@@ C.1 Derivation of the Linearized QRE Constraints
$\frac{\mu^{*}(a)}{\mu^{*}(1)}=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)/Z}{\exp(Q(1,\cdot)\nu^{*}/\tau)/Z}=\frac{\exp(Q(a,\cdot)\nu^{*}/\tau)}{\exp(Q(1,\cdot)\nu^{*}/\tau)},$ (C.2)
@@
$\log\left(\frac{\mu^{*}(a)}{\mu^{*}(1)}\right)$ $=\log(\exp(Q(a,\cdot)\nu^{*}/\tau))-\log(\exp(Q(1,\cdot)\nu^{*}/\tau))$ (C.3)
$=\frac{1}{\tau}(Q(a,\cdot)\nu^{*}-Q(1,\cdot)\nu^{*})$ (C.4)
$=\frac{1}{\tau}(Q(a,\cdot)-Q(1,\cdot))\nu^{*}.$ (C.5)
Rearranging the terms yields the first set of linear constraints in Eq. 3.4:
$(Q(a,\cdot)-Q(1,\cdot))\nu^{*}=\tau\cdot\log\left(\frac{\mu^{*}(a)}{\mu^{*}(1)}\right),\quad\forall a\in\mathcal{A}\setminus\{1\}.$ (C.6)
Similarly, for Player 2 (Eq. 3.3), we analyze the ratio of $\nu^{*}(b)$ to $\nu^{*}(1)$:
$\frac{\nu^{*}(b)}{\nu^{*}(1)}=\frac{\exp(-Q(\cdot,b)^{\top}\mu^{*}/\tau)}{\exp(-Q(\cdot,1)^{\top}\mu^{*}/\tau)}.$ (C.7)
@@
$\log\left(\frac{\nu^{*}(b)}{\nu^{*}(1)}\right)$ $=\frac{1}{\tau}(-Q(\cdot,b)^{\top}\mu^{*}-(-Q(\cdot,1)^{\top}\mu^{*}))$ (C.8)
$=\frac{1}{\tau}(Q(\cdot,1)^{\top}-Q(\cdot,b)^{\top})\mu^{*}$ (C.9)
$=\frac{1}{\tau}(Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}.$ (C.10)
Rearranging the terms yields the second set of linear constraints in Eq. 3.4:
$(Q(\cdot,1)-Q(\cdot,b))^{\top}\mu^{*}=\tau\cdot\log\left(\frac{\nu^{*}(b)}{\nu^{*}(1)}\right),\quad\forall b\in\mathcal{B}\setminus\{1\}.$ (C.11)
@@
C.2 Key Definitions and Theorems from Matrix Analysis
@@
Definition C.1 (Matrix Norms).
@@
Definition C.2 (Moore-Penrose Pseudoinverse).
@@
Theorem C.1 (Weyl’s Inequality for Singular Values).
@@
This theorem is crucial for establishing the stability of the matrix $X$ under small perturbations (Step 1 in the proof of Theorem 4.6).
C.3 Key Theorems from Concentration Inequalities
@@
Theorem C.2 (Bretagnolle-Huber-Carol Inequality).
@@
$P(\|\hat{\mu}-\mu^{*}\|_{1}>\epsilon)\leq 2^{m}e^{-N\epsilon^{2}/2}.$ (C.13)
This theorem provides the finite-sample bound on the L1 distance between the estimated and true policies.
Appendix D Proofs for Matrix Games
@@
D.1 Proof of Theorem 4.4 (Identifiability of Blind-IGT)
@@
Proof of Sufficiency ($\Rightarrow$): Assume the Rank Condition ($\text{rank}(X^{*})=d$) and the Non-Uniformity Condition ($y^{*}\neq 0$) hold.
@@
$\theta^{*}=(X^{*})^{\dagger}(\tau^{*}y^{*})=\tau^{*}(X^{*})^{\dagger}y^{*}.$ (D.1)
@@
$\|\theta^{*}\|_{2}=\|\tau^{*}\theta_{LS}^{*}\|_{2}=|\tau^{*}|\|\theta_{LS}^{*}\|_{2}=1.$ (D.2)
@@
$\theta^{*}=\tau^{*}\theta_{LS}^{*}=\frac{\theta_{LS}^{*}}{\|\theta_{LS}^{*}\|_{2}}.$ (D.4)
Proof of Necessity ($\Leftarrow$): We prove the contrapositive by considering the failure of each condition.
@@
D.2 Supporting Lemmas for Finite Sample Analysis
We present and prove the supporting lemmas required for the proof of Theorem 4.6. These lemmas characterize the concentration of the policy estimates and the resulting perturbations in the system matrices.
Lemma D.1 (Policy Estimation Error).
@@ D.2 Supporting Lemmas for Finite Sample Analysis
By the Bretagnolle-Huber-Carol inequality (Theorem C.2 in Appendix C), for any $\epsilon>0$:
$P(\|\hat{\mu}-\mu^{*}\|_{1}>\epsilon)\leq 2^{m}e^{-N\epsilon^{2}/2}.$ (D.5)
@@
$\frac{\delta}{2}=2^{m}e^{-N\epsilon_{\mu}^{2}/2}\implies\log(2\cdot 2^{m}/\delta)=N\epsilon_{\mu}^{2}/2\implies\epsilon_{\mu}=\sqrt{\frac{2\log(2\cdot 2^{m}/\delta)}{N}}.$ (D.6)
@@
Lemma D.2 (Perturbation of X and y).
Let $L$ be the bound on $\|\phi\|_{2}$ (Assumption 4.1). Under Assumption 4.5, let $\xi>0$ be the minimum probability. If $N$ is large enough such that the policy estimation errors satisfy $\epsilon_{\mu}<\xi/2$ and $\epsilon_{\nu}<\xi/2$, then with high probability (conditioned on the event of Lemma D.1):
@@
$\|(A(\hat{\nu})-A(\nu^{*}))_{a-1}\|_{2}$ $=\left\|\sum_{b^{\prime}\in\mathcal{B}}(\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime}))(\phi(a,b^{\prime})-\phi(1,b^{\prime}))^{\top}\right\|_{2}$ (D.9)
$\leq\sum_{b^{\prime}\in\mathcal{B}}|\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime})|\cdot\|\phi(a,b^{\prime})-\phi(1,b^{\prime})\|_{2}.$ (D.10)
@@
$\|(A(\hat{\nu})-A(\nu^{*}))_{a-1}\|_{2}$ $\leq 2L\sum_{b^{\prime}\in\mathcal{B}}|\hat{\nu}(b^{\prime})-\nu^{*}(b^{\prime})|=2L\|\hat{\nu}-\nu^{*}\|_{1}\leq 2L\epsilon_{\nu}.$ (D.11)
@@
$\|A(\hat{\nu})-A(\nu^{*})\|_{F}\leq\sqrt{(m-1)(2L\epsilon_{\nu})^{2}}\leq 2L\sqrt{m}\epsilon_{\nu}.$ (D.12)
@@
$\|\hat{X}-X^{*}\|_{op}$ $\leq\|\hat{X}-X^{*}\|_{F}=\sqrt{\|A(\hat{\nu})-A(\nu^{*})\|_{F}^{2}+\|B(\hat{\mu})-B(\mu^{*})\|_{F}^{2}}$ (D.13)
$\leq\sqrt{4L^{2}m\epsilon_{\nu}^{2}+4L^{2}n\epsilon_{\mu}^{2}}=2L\sqrt{m\epsilon_{\nu}^{2}+n\epsilon_{\mu}^{2}}.$ (D.14)
@@
$\|c(\hat{\mu})-c(\mu^{*})\|_{2}^{2}$ $=\sum_{a=2}^{m}|c(\hat{\mu})_{a-1}-c(\mu^{*})_{a-1}|^{2}$ (D.20)
$\leq\sum_{a=2}^{m}\left(\frac{2}{\xi}\right)^{2}(|\hat{\mu}(a)-\mu^{*}(a)|+|\hat{\mu}(1)-\mu^{*}(1)|)^{2}$ (D.21)
$\leq\frac{8}{\xi^{2}}\sum_{a=2}^{m}(|\hat{\mu}(a)-\mu^{*}(a)|^{2}+|\hat{\mu}(1)-\mu^{*}(1)|^{2})\quad(\text{Since }(x+y)^{2}\leq 2(x^{2}+y^{2}))$ (D.22)
$=\frac{8}{\xi^{2}}\left(\sum_{a=2}^{m}|\hat{\mu}(a)-\mu^{*}(a)|^{2}+(m-1)|\hat{\mu}(1)-\mu^{*}(1)|^{2}\right)$ (D.23)
$\leq\frac{8m}{\xi^{2}}\|\hat{\mu}-\mu^{*}\|_{2}^{2}.$ (D.24)
@@
$\|c(\hat{\mu})-c(\mu^{*})\|_{2}\leq\frac{\sqrt{8m}}{\xi}\epsilon_{\mu}.$ (D.25)
@@
$\|\hat{y}-y^{*}\|_{2}$ $=\sqrt{\|c(\hat{\mu})-c(\mu^{*})\|_{2}^{2}+\|d(\hat{\nu})-d(\nu^{*})\|_{2}^{2}}$ (D.26)
$\leq\sqrt{\frac{8m}{\xi^{2}}\epsilon_{\mu}^{2}+\frac{8n}{\xi^{2}}\epsilon_{\nu}^{2}}=\frac{\sqrt{8}}{\xi}\sqrt{m\epsilon_{\mu}^{2}+n\epsilon_{\nu}^{2}}.$ (D.27)
@@
Lemma D.3 (Stability of Normalization).
@@
$\left\|\frac{v_{1}}{\|v_{1}\|_{2}}-\frac{v_{0}}{\|v_{0}\|_{2}}\right\|_{2}=\left\|\frac{v_{1}\|v_{0}\|_{2}-v_{0}\|v_{1}\|_{2}}{\|v_{1}\|_{2}\|v_{0}\|_{2}}\right\|_{2}.$ (D.28)
@@
$\|v_{1}\|v_{0}\|_{2}-v_{0}\|v_{1}\|_{2}\|$ $=\|(v_{0}+\Delta v)\|v_{0}\|_{2}-v_{0}(\|v_{0}\|_{2}+\Delta_{\|v\|})\|$ (D.29)
$=\|v_{0}\|v_{0}\|_{2}+\Delta v\|v_{0}\|_{2}-v_{0}\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|$ (D.30)
$=\|\Delta v\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|.$ (D.31)
@@
$\|\Delta v\|v_{0}\|_{2}-v_{0}\Delta_{\|v\|}\|$ $\leq\|\Delta v\|v_{0}\|_{2}\|+\|v_{0}\Delta_{\|v\|}\|$ (D.32)
$=\|\Delta v\|_{2}\|v_{0}\|_{2}+|\Delta_{\|v\|}|\|v_{0}\|_{2}$ (D.33)
$\leq\epsilon\|v_{0}\|_{2}+\epsilon\|v_{0}\|_{2}=2\epsilon\|v_{0}\|_{2}.$ (D.34)
@@
$\left\|\frac{v_{1}}{\|v_{1}\|_{2}}-\frac{v_{0}}{\|v_{0}\|_{2}}\right\|_{2}\leq\frac{2\epsilon\|v_{0}\|_{2}}{(\delta/2)\|v_{0}\|_{2}}=\frac{4\epsilon}{\delta}.$ (D.36)
D.3 Proof of Theorem 4.6 (Finite Sample Bounds)
@@
Let $\sigma^{*}=\sigma_{\min}(X^{*})$. Since the Rank Condition holds (Theorem 4.4), $\sigma^{*}>0$. Let $\epsilon_{N}$ be the policy estimation error rate from Lemma D.1. Let $C_{X},C_{Y}$ be the perturbation constants from Lemma D.2.
@@
Stability of the Estimated System. We need to ensure that the estimated matrix $\hat{X}$ remains full rank and well-conditioned, provided $N$ is sufficiently large. We use Weyl’s inequality (Theorem C.1 in Appendix C) for the perturbation of singular values:
$\sigma_{\min}(\hat{X})\geq\sigma_{\min}(X^{*})-\|\hat{X}-X^{*}\|_{op}=\sigma^{*}-\|\hat{X}-X^{*}\|_{op}.$ (D.37)
We require $\|\hat{X}-X^{*}\|_{op}<\sigma^{*}/2$ to guarantee $\sigma_{\min}(\hat{X})>\sigma^{*}/2$. By Lemma D.2, this holds if $C_{X}\epsilon_{N}<\sigma^{*}/2$. This defines the primary condition for "sufficiently large NN". We also require $N$ large enough such that the conditions of Lemma D.2 (i.e., $\epsilon_{\mu},\epsilon_{\nu}<\xi/2$) hold.
@@
We utilize standard perturbation bounds for the least squares solution when the true system is consistent (zero residual). Since $\hat{X}$ is full rank (Step 1), $\hat{\theta}_{LS}=\hat{X}^{\dagger}\hat{y}$. The error bound is given by (e.g., Theorem 20.1 in Higham [2002], specialized for zero residual):
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}\leq\|\hat{X}^{\dagger}\|_{2}(\|\hat{X}-X^{*}\|_{op}\|\theta^{\prime*}\|_{2}+\|\hat{y}-y^{*}\|_{2})+\mathcal{O}(\epsilon_{N}^{2}).$ (D.38)
Since $\sigma_{\min}(\hat{X})\geq\sigma^{*}/2$, we have $\|\hat{X}^{\dagger}\|_{2}=1/\sigma_{\min}(\hat{X})\leq 2/\sigma^{*}$. We substitute the bounds from Lemma D.2 and note that $\|\theta^{\prime*}\|_{2}=\|\theta^{*}\|_{2}/\tau^{*}=1/\tau^{*}$ (since $C=1$).
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}$ $\leq\frac{2}{\sigma^{*}}\left((C_{X}\epsilon_{N})\frac{1}{\tau^{*}}+C_{Y}\epsilon_{N}\right)$ (D.39)
$=\left(\frac{2C_{X}}{\sigma^{*}\tau^{*}}+\frac{2C_{Y}}{\sigma^{*}}\right)\epsilon_{N}=:C_{LS}\epsilon_{N}.$ (D.40)
@@
Error in $\hat{\theta}$. We apply Lemma D.3 (Stability of Normalization) with $v_{1}=\hat{\theta}_{LS}$ and $v_{0}=\theta^{\prime*}$. The error is $\epsilon=\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}\leq C_{LS}\epsilon_{N}$. The norm is $\delta=\|\theta^{\prime*}\|_{2}=1/\tau^{*}$. We require $\epsilon<\delta/2$, i.e., $C_{LS}\epsilon_{N}<1/(2\tau^{*})$. This holds for sufficiently large $N$ since $\epsilon_{N}\rightarrow 0$.